OpenStudy (anonymous):

When the volume of a cell increases, its surface area does what? I know that the surface area increases but does it increase at a faster rate, slowere rate, or at the same rate?

OpenStudy (anonymous):

at the same rate I think

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

volume increases with the radius cubed, surface area increases with the radius squared. V = (4/3)(pi)(r)^3 SA = (4)(pi)(r)^2